DGD Bond Practice Questions: 1. Apple issues an 8 year bond with a par value of S1050. The bond has a 6% annual coupon rate and a face value of $1000 a) What is the bonds YTM (Yield to maturity)? b) If the bond can be called in 4 years, what is the YTC (Yield to call)? The price dropped to 1025
Answer #1

Answer A) Maturity years = 8 years

Face Value = $1000

Annual Coupon rate = 6% = 1000*6% = $ 60

Price = $ 1050

YTM ( Yield to maturity ) = 60+(1000-1050)/8/(1000+1050)/2

=(60-6.25)/1025

=2.2439%

b) Coupon payment for annual fixed payment= 60

Call price = 1025

Number of years to call = 4 years

Face vlue of bond = 1000

Yield to call (YTC) = 60+(1000-1025)/4/(1000+1025)/2

= (60-6.25)/1012.5

= 5.3086%